OpenStudy (anonymous):

on a trip a family averaged 65km/hr for the first hour 50km/hr for the second hour and 58 km/hr for the third hour. what was their average speed?

mathslover (mathslover):

a physics question @alfers101 please post this in physics section

mathslover (mathslover):

though I will help you wait

mathslover (mathslover):

given : 1) at t =1 hr. , average speed = 65 kmph 2) at t = 2hr. , average speed = 58 kmph so we know that: \[\large{\frac{\textbf{total distance travelled}}{\textbf{total time taken}}=\textbf{average speed}}\]

mathslover (mathslover):

Oh sorry wait

OpenStudy (lgbasallote):

also remember that \[\huge \frac{\text{speed}_1 + \text{speed}_2}{2} \ne \text{average speed}\] just saying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

mathslover (mathslover):

thanks? u got it soo soon @alfers101 ?

OpenStudy (anonymous):

no i still dont :( but i have the idea but i dont know how to answer this

mathslover (mathslover):

oh k so what is the idea?

mathslover (mathslover):

do you have the answer alfers101?

OpenStudy (phi):

FIrst, find how far they drove. Use speed*time= distance they drove 65km/hr for the first hour: distance= 65*1 do the same thing for the 2nd hour (using the new speed), and the 3rd hour add up the distances now use speed*time= distance speed= distance/time distance will be the total distance time will be the total time ( 3 hours)